# Hadamard Representations: Augmenting Hyperbolic Tangents in RL

### v3

@@ A.1 Hyperparameters
Hadamard Implementation
Constructing a Hadamard representation is a straightforward process that only requires an additional, parallel linear layer. Starting from the flattened convolutional features in DQN, the Pytorch pseudocode is defined as follows:
flattened_features = conv_features.flatten(1)
representation1 = nn.Tanh(linear1(flattened_features))
representation2 = nn.Tanh(linear2(flattened_features))
hadamard = representation1 * representation2
Q-values = linear_q(hadamard)
Reinforcement Learning
@@
A.2 PPO Architecture
Figure 11: A visualisation of the Hadamard representation (HR) architecture combined with the PPO architecture applied on a snapshot of the ’SpaceInvaders’ Atari environment. A parallel linear layer providing an independent representation is integrated, where the Hadamard product of the two parallel representations represents the final representation $z_{t}$.
@@ Appendix B Kernel Density Estimations
As discussed in Section 3, we hypothesize that the differences between a hyperbolic tangent with and without an HR are due to the increased ability of the product of hyperbolic tangents being able to negate dying neurons. We further see this phenomenom when plotting a random selection of neurons from both the mask and the base representation in Fig. 12(a).
@@ C.4 ReLU activated Hadamard representation
C.5 Hyperbolic Tangent Convolutional Activation
We additionally run tests in DQN on 8 Atari games. The baseline ReLU is used, except for the convolutional activations which are changed to a hyperbolic tangent.
Figure 18: Baseline normalized performance in the Atari domain for 10M iterations (40M Frames) on DQN.
It is evident that using non-Hadamard hyperbolic tangents in the convolutional layers can lead to only a fraction of the performance of the ReLU-activated baseline.
@@ Appendix D Atari
D.1 DQN Evaluation Details
We normalize performance with respect to the ReLU baseline in cleanrl  Huang et al. (2022). The minimum and maximum score of the ReLU baseline are taken for each environment, and the normalized score for each environment is calculated as follows:
@@ D.2 Individual Environment Scores
Figure 20: DQN Performance comparison on the individual Atari Environments. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region. Eight popular Atari games in the RL community that are non-exploration driven were chosen due to computational limitations. No other games were tested.
Figure 21: PPO Performance comparison on the individual Atari Environments. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region. Eight popular Atari games in the RL community that are non-exploration driven were chosen due to computational limitations. No other games were tested.
Figure 22: PQN Performance comparison on the individual Atari Environments. Labels represent MLP / CNN activations. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
Figure 23: PQN Performance comparison on the individual Atari Environments. Labels represent MLP / CNN activations. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
Figure 24: PQN Performance comparison on the individual Atari Environments. Labels represent MLP / CNN activations. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
Figure 25: PQN Performance comparison on the individual Atari Environments. Labels represent MLP / CNN activations. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
Figure 26: PQN Performance comparison on the individual Atari Environments. Labels represent MLP / CNN activations. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
Figure 27: PQN Performance comparison on the individual Atari Environments. Labels represent MLP / CNN activations. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
Figure 28: PQN Performance comparison on the individual Atari Environments. Labels represent MLP / CNN activations. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.

### v4

@@ A.1 Hyperparameters
A.2 Hadamard Implementation
Constructing a Hadamard representation is a straightforward process that only requires additional, parallel incoming weights. Starting from a hidden layer, the Pytorch pseudocode is defined as follows:
# hidden = previous hidden layer
linear1 = nn.Linear(input_dim, output_dim)
linear2 = nn.Linear(input_dim, output_dim)
representation1 = nn.Tanh(linear1(hidden))
representation2 = nn.Tanh(linear2(hidden))
hadamard_representation = representation1 * representation2
A.3 Reinforcement Learning
@@ Appendix B Kernel Density Estimations
As discussed in Section 4, we hypothesize that the differences between a hyperbolic tangent with and without an HR are due to the increased ability of the product of hyperbolic tangents being able to negate dying neurons. We further see this phenomenon when plotting a random selection of neurons from both the mask and the base representation in Fig. 10(a).
@@ Appendix D Atari
D.1 DQN & PPO Evaluation Details
For DQN and PPO, the Hadamard representation was applied to the final hidden layer of the Nature CNN. Furthermore, we have normalized performance with respect to the ReLU baseline on which the experiments were build (Huang et al., 2022). The minimum and maximum score of the ReLU baseline are taken for each environment, and the normalized score for each environment is calculated as follows:
@@ D.2 Individual Environment Scores
Figure 17: DQN Performance comparison on the individual Atari Environments. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
Figure 18: PPO Performance comparison on the individual Atari Environments. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
Figure 19: PQN Performance comparison on the individual Atari Environments. Labels represent encoder activations. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
Figure 20: PQN Performance comparison on the individual Atari Environments. Labels represent encoder activations. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
Figure 21: PQN Performance comparison on the individual Atari Environments. Labels represent encoder activations. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
Figure 22: PQN Performance comparison on the individual Atari Environments. Labels represent encoder activations. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
Figure 23: PQN Performance comparison on the individual Atari Environments. Labels represent encoder activations. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
Figure 24: PQN Performance comparison on the individual Atari Environments. Labels represent encoder activations. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.
Figure 25: PQN Performance comparison on the individual Atari Environments. Labels represent encoder activations. Plotted lines represent the mean taken over 5 seeds, with the standard deviations expressed as the shaded region.